### v1

# SceneSplat: Gaussian Splatting-based Scene Understanding with Vision-Language Pretraining

Abstract
Recognizing arbitrary or previously unseen categories is essential for comprehensive real-world 3D scene understanding. Currently, all existing methods rely on 2D or textual modalities during training, or together at inference. This highlights a clear absence of a model capable of processing 3D data alone for learning semantics end-to-end, along with the necessary data to train such a model. Meanwhile, 3D Gaussian Splatting (3DGS) has emerged as the de facto standard for 3D scene representation across various vision tasks. However, effectively integrating semantic reasoning into 3DGS in a generalizable fashion remains an open challenge. To address these limitations we introduce SceneSplat in Fig. 1, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3DGS. Furthermore, we propose a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. In order to power the proposed methods, we introduce SceneSplat-7K, the first large-scale 3DGS dataset for indoor scenes, comprising of 6868 scenes derived from 7 established datasets like ScanNet, Matterport3D, etc. Generating SceneSplat-7K required computational resources equivalent to 119 GPU-days on an L4 GPU, enabling standardized benchmarking for 3DGS-based reasoning for indoor scenes. Our exhaustive experiments on SceneSplat-7K demonstrate the significant benefit of the proposed methods over the established baselines. Our code, model, and datasets will be released at SceneSplat.
Figure 1: We present the 3DGS indoor dataset SceneSplat-7K which includes 7K scenes generated from ARKitScenes [1], Replica [45], ScanNet [5], ScanNet++ [57], Hypersim[42], 3RScan [48], and Matterport3D [2]. Leveraging this high-quality dataset, we propose SceneSplat, the first model to predict open-vocabulary language features for millions of 3D Gaussians in a single forward pass.
$\ast$ indicates equal contribution. $\dagger$ indicates the corresponding author: Bin Ren <bin.ren@insait.ai>.
1 Introduction
The ability to interpret arbitrary queries, rather than being limited to a closed-set of categories, is crucial for 3D understanding models to generalize across diverse real-world scenarios. Traditional 3D vision systems are typically trained on fixed closed-set category labels drawn from datasets like ScanNet [5]. Such label spaces fail to capture the diversity of concepts in real environments. This gap has spurred significant interest in open-vocabulary 3D scene understanding, which aims to recognize arbitrary or unseen categories beyond a predefined taxonomy [8, 54, 15]. Achieving this capability would empower the model to reason about novel objects in a scene via natural language descriptors.
Achieving open-vocabulary recognition in 3D is challenging due to the scarcity of large-scale 3D-text paired data. Breakthroughs in 2D vision, driven by internet-scale image-text pre-training, cannot be directly leveraged in 3D due to the absence of analogous 3D datasets with rich textual annotations. To address this gap, current methods resort to multi-modality fusion, distilling knowledge from 2D vision-language models into 3D data. Some approaches project 3D points into source images and use a pretrained 2D backbone (e.g., CLIP) to supervise 3D feature learning [64, 33, 70]; others generate synthetic captions for 3D scenes to explicitly associate point clouds with semantic-rich text [43, 7]. However, all current methods rely on 2D or textual modalities during training, or together at inference, to compensate for limited 3D semantics. This highlights a key limitation: the absence of a robust model for processing 3D data end-to-end for semantic learning, along with the lack of sufficient data for training such a model.
The field of 3D representation is rapidly evolving. While classical 3D networks rely on point clouds or voxel grids [55, 50, 51], recent approaches inspired by radiance fields have greatly improved scene representation. Notably, 3D Gaussian Splatting (3DGS) [16] has emerged as an efficient representation, achieving state-of-the-art performance in view synthesis and geometry modeling [58, 61]. Unlike discrete point clouds, 3DGS provides a compact formulation that can be optimized per scene for 2D synthesis, while the underlying Gaussians also softly encode rich 3D structures (positions, shapes, and opacities of the regions). This makes it a unique candidate for 3D scene cues, as it naturally fuses geometry and appearance information. However, integrating semantic reasoning into 3DGS is non-trivial. The naïve way of optimizing additional semantic features in 3DGS [17] is inefficient and limited to a single scene. Consequently, generalizable open-vocabulary understanding in 3DGS remains unexplored, as no existing model directly processes 3D Gaussian parameters. Moreover, no large-scale scene-level 3DGS dataset is available to train such models, further limiting progress in this direction.
We address these limitations with SceneSplat, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3D Gaussian splats. The proposed model is powered by SceneSplat-7K, a carefully curated 3DGS indoor scene dataset spanning around 7K scenes. SceneSplat introduces a 3DGS encoder that takes as input the parameters of a Gaussian-splat scene (center, scale, color, opacity) and outputs semantic features in a per-primitive manner, in a single forward pass. We leverage supervision from vision-language models to train SceneSplat’s encoder to produce CLIP-aligned embeddings, effectively bridging language and 3D without an explicit 2D fusion at runtime. Furthermore, we propose GaussSSL, a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. GaussSSL operates through three synergistic strategies, Masked Gaussian Modeling (MGM), Self-Distillation, and optionally Language-Gaussian Feature Alignment, allowing the model to separate high-level semantic signals from raw parameters alone. This self-supervised pre-training is fueled by large amounts of 3DGS scenes in the SceneSplat-7K dataset. Our contributions can be summarized as follows:
$\bullet$ We present SceneSplat-7K, a high-quality large-scale Gaussian splats dataset spanning 7K indoor scenes, which boost 3DGS scene understanding research.
$\bullet$ We propose SceneSplat, the model that unlocks open-vocabulary recognition for 3D Gaussian splats and achieves state-of-the-art zero-shot semantic segmentation performance on three fine-grained benchmarks.
$\bullet$ We incorporate annotation-free self-supervised training mechanisms on this large-scale indoor dataset and demonstrate their effectiveness in downstream indoor segmentation task.
2 Related Work
Metric ScanNet[5] ScanNet++[57] ScanNet++v2[57] Replica[45] Hypersim[42] 3RScan[48] ARKitScenes[1] Matterport3D[2] Scenesplat-7K
Raw Scenes 1513 380 1006 18 461 1482(scans) 1661 2194 8715 $\sim$9k
GS Scenes 1513 281 956 18 448 632(scans) 938 1982 6868 $\sim$7k
RGB Frames 2.5M 228K 1.1M 36K 77K 156K 450K 194K 4.74M
Storage 1.1TB 128GB 95GB 14GB 236GB 439GB 159GB 95GB 2.27TB
PSNR 29.07 29.49 29.11 41.25 25.93 27.46 25.65 28.64 28.17
Depth Loss 0.031 0.019 0.015 0.002 0.228 0.018 0.0136 0.122 0.064
SSIM 0.869 0.924 0.933 0.980 0.894 0.881 0.841 0.900 0.889
LPIPS 0.236 0.133 0.116 0.0396 0.157 0.335 0.380 0.208 0.230
GS per scene 1.50M 1.89M 1.89M 1.50M 2.84M 1.50M 1.52M 0.50M 1.37M
Total GS 2,269.5M 528.9M 1810.3M 23.95M 1,237.5M 948.0M 343.9M 991.0M 8.15B
GPU Time (L4) 593h 177h 594h 8h 176h 576h 290h 443h 2857h
Table 1: Dataset Statistics. The proposed SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from Scannet[5], Scannet++[57], Scannet++v2, Replica[45], Hypersim[42], 3RScan[48], arkitscenes[1], and matterport3d[2]. The dataset contains 6,868 scenes and 8.15 Billion 3DGS. Constructing this dataset required computational resources equivalent to 119 GPU-days on one NVIDIA L4 GPU. SceneSplat-7K achieves high-fidelity reconstruction quality with an average PSNR 28.17 dB.
3D Indoor Datasets. The advancement of 3D deep learning has driven the development of various indoor datasets for scene understanding, reconstruction, and representation learning [5, 57, 42, 1, 45, 48, 2]. ScanNet [5] serves as a fundamental benchmark with 1,513 real-world indoor scenes and dense annotations, while ScanNet++ [57] extends it with additional real and synthetic scenes. Hypersim [42] offers photorealistic synthetic environments, and ARKitScenes [1] captures 1,661 real-world indoor scenes using ARKit. Replica [45] provides high-fidelity reconstructions, and 3RScan [48] supports multi-view analysis with 1,482 scans. Habitat-Matterport3D [2] integrates Matterport3D into an embodied AI simulation platform. These datasets are essential for 3D perception but lack large-scale support for emerging 3D representations like 3DGS. Prior works [13, 29] have explored NeRF-based representations, and ShapeSplat [28] introduced Gaussian-splatted objects, but no dataset currently supports indoor scene understanding with Gaussian Splatting. To address this limitation, we introduce a 3D Gaussian Splatting scene dataset derived from 7 widely used indoor datasets[5, 57, 42, 1, 45, 48, 2]. Our dataset offers a rich collection of indoor scenes from both real-world and synthetic sources, featuring high-quality Gaussian splats, facilitating the transition from object-level to scene-level Gaussian splatting and advancing large-scale 3D scene understanding.
Open Vocabulary Scene Understanding. Recent advancements in open-vocabulary models have significantly expanded the capabilities of vision-language understanding. Foundation models such as DINO [62, 31] have enabled self-supervised visual feature extraction at scale, effectively supporting tasks like detection and segmentation. SAM [20, 39] introduced the capability of prompt-driven segmentation, generalizing robustly across diverse datasets and tasks. CLIP [37] pioneered aligning visual and textual embedding spaces, allowing for zero-shot transfer across numerous downstream tasks. Subsequently, SigLIP [60, 46] improved alignment by introducing non-linear activation techniques, enhancing open-vocabulary performance marginally. Recent works [69, 67, 36, 4, 33, 10] have also introduced 2D open-vocabulary foundation models into 3D by utilizing NeRF [30] or 3DGS [16]. LERF [17, 38] integrated language queries into NeRF-based models, enabling rich semantic querying within 3D scenes. LangSplat [36] combined 3DGS with open-vocabulary embeddings from SAM [39] and CLIP [37], facilitating open-vocabulary semantic scene understanding. OccamLGS [4] further optimized this process by employing feature lifting, directly projecting all 2D features into 3DGS, and reducing computation time from hours to seconds. However, these methods require time-consuming preprocessing of images using 2D foundation models. In contrast, we propose a pipeline and dataset for training a 3D foundation model, enabling feed-forward open-vocabulary understanding of 3DGS.
3D Representation Learning. Representation learning in 3D, akin to its success in the 2D image domain, is crucial for extracting meaningful features from 3D data. Previous approaches have focused on either architectural advancements or learning paradigms. Many methods have sought to map 3D data into 2D patches, leveraging conventional 2D CNNs or Vision Transformers [47, 9, 40]. Although promising, this approach often overlooks the inherent properties of 3D data due to the embedding strategy. To address this, research has shifted toward specialized architectures (i.e., PointCNN [22], PointNet [34], and Point Transformer [65, 50, 51], Pointmamba [24]). However, labeling 3D data requires extensive effort and domain expertise, making it quite challenging. This has highlighted the importance of self-supervised learning (SSL) for 3D data, with methods typically falling into contrastive or generative categories. Contrastive methods aim to differentiate similar and dissimilar examples, but often suffer from issues such as overfitting and mode collapse, exacerbated by the sparse nature of 3D data [35, 41]. On the other hand, generative methods, inspired by BERT’s mask-and-reconstruct strategy [6], have proven more effective. Masked Autoencoders [11], originally designed for 2D images, have been adapted for 3D data types, including point clouds [32, 63], meshes [25], voxels [12], and Gaussians [28], allowing the reconstruction of masked regions and facilitating the learning of robust spatial and semantic features. In this work, we propose a novel approach for language-aware 3DGS pretraining, as well as a self-supervised learning scheme that operates via Gaussian masking and self-distillation.
3 SceneSplat Dataset
We introduce SceneSplat-7K – a carefully curated dataset of 3D Gaussian Splats representing indoor scenes. The main goal was obtaining a dataset to facilitate generalizable 3DGS indoor scene understanding. The dataset contains about seven thousand scenes, including both real-world and synthetic environments. The comprehensive statistics of the introduced dataset are presented in Tab. 1. The SceneSplat-7K dataset will be publicly released, with additional details (license) available in the supplement.
3.1 Data Processing
Multiple measures are applied before, during, and after the 3DGS optimization to guarantee high-quality 3DGS scenes. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. We remove blurry frames by using the variance of the Laplacian as a sharpness metric. We use gsplat [56] for 3DGS optimization. For scenes with available depth input, we apply depth loss to achieve better geometry modeling. To efficiently compress the 3DGS scene, we employ a Markov Chain Monte Carlo strategy [18] and add opacity and scale regularization. Once optimized, we filter 3DGS scenes based on the PSNR metric before using them as input for our pretraining. We refer to the supplement for per-dataset processing details.
3.2 Data Statistic
SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from Scannet[5], Scannet++[57], Scannet++v2, Replica[45], Hypersim[42], 3RScan[48], Arkitscenes[1], and Matterport3D[2], comprising approximately 9,000 raw scenes. Scenesplat-7K contains 6,868 processed Gaussian splatting scenes, with an average of 1.37 Million 3D Gaussians per scene and 8.15 Billion Gaussians in total, from which 4,114 high-quality Gaussian splatting scenes are selected for pertaining. Spanning a total of 4.74 Million RGB frames, SceneSplat-7K achieves high-fidelity appearance and competitive reconstruction quality, having an average PSNR of 28.17 dB, average depth loss of 0.064 m, average SSIM of 0.889, and average LPIPS of 0.230. Constructing this dataset required the equivalent of 119 days of computation on an NVIDIA L4 GPU.
Figure 2: SceneSplat Overview. The SceneSplat-7K dataset enables Vision-Language Pretraining and Self-Supervised Pretraining. For vision-language pretraining, we associate each 3D Gaussian primitive with semantic features based on our label collection process and train a generalizable open-vocabulary learner that predict per-gaussian embeddings. For self-supervised pretraining, we employ Masked Gaussian Modeling to reconstruct masked primitives, Self-Distillation Learning for augmentation-invariant features, and Language-Gaussian Alignment for scenes with collected labels. The former achieves state-of-the-art zero-shot segmentation results on ScanNet200 [5], ScanNet++ [57], and Matterport3D [2] benchmarks and the latter unlocks training on large-scale 3DGS data.
4 Methodology
Building upon the SceneSplat-7K dataset, we carry out both vision-language 3DGS pretraining which enables open-vocabulary scene understanding, and self-supervised pretraining which regularizes the latent space during 3DGS parameter encoding as shown in Fig. 2. For vision-language pretraining, we first need to collect the primitive-level language labels for the 3DGS scenes (Sec. 4.1). The SceneSplat model then learns to predict vision-language features robustly from Gaussian parameters and its surroundings (Sec. 4.2). Furthermore, for self-supervised pretraining, we employ a multi-objective self-supervised training framework, integrating reconstruction and self-distillation alignment (Sec. 4.3).
4.1 3DGS Language Label Collection
Our language label collection aims to establish 3D-language paired data by associating each 3D Gaussian primitive $G_{i}$ with a rich semantic feature $F_{i}\in\mathbb{R}^{d}$.
Unlike methods that align 3D primitives with text embeddings [8, 21] or use visual captioning [26, 59], we directly align Gaussians with the image embedding space of vision-language models (VLM), preserving richer latent semantic information. Our approach also avoids scene-specific compression methods [36, 44] that limit scalability and generalization.
As outlined in Alg. 1, we employ SAMv2 [39] for object-level segmentation and SigLIP2 [46] for feature extraction. We then use Occam’s LGS [4] to efficiently lift these 2D feature maps to a 3D Gaussian feature field in a optimization-free manner. This results in a comprehensive collection of 3DGS-feature pairs $\{(G_{i},F_{i})\}_{i=1}^{N}$ across multiple datasets, providing a solid foundation for our vision-language pretraining.
Algorithm 1 3DGS Language Label Collection
1:Input: Training views $\{I_{j}\}_{j=1}^{M}$, 3D Gaussian scene $\{G_{i}\}_{i=1}^{N}$, SAMv2, SigLip2
2:Output: 3D Gaussian-feature pairs $\{(G_{i},F_{i})\}_{i=1}^{N}$
3:Step 1: 2D Feature Map Generation
4:for each training view $I_{j}$ do
5:     $M_{\text{seg}}\leftarrow\text{SAMv2}(I_{j})$triangleright Get object-level seg. masks
6:     $f_{g}\leftarrow\text{SigLip2}(I_{j})$triangleright Feature from full frame
7:     Initialize feature map $F_{j}$ for view $I_{j}$
8:     for each segment $s$ in $M_{\text{seg}}$ do
9:         $f_{l},f_{m}\leftarrow\text{SigLip2}(\text{crop}(I_{j},s))$
10$\triangleright$ Local features for crops w/ and w/o background
11:         Dynamic Weighting:
12:         $w_{g},w_{l},w_{m}\leftarrow\text{compute\_weights}(f_{g},f_{l},f_{m})$
13$\triangleright$ Weights based on context
14:         $f_{s}\leftarrow w_{g}\cdot f_{g}+w_{l}\cdot f_{l}+w_{m}\cdot f_{m}$
15$\triangleright$ Fuse features
16:         Update feature map $F_{j}$ with $f_{s}$ at segment $s$
17:     end for
18:end for
19:Step 2: Lifting 2D Features to 3D Gaussian Feature Field
20:$\{F_{i}\}_{i=1}^{N}\leftarrow\text{Occam's\_LGS}(\{F_{j}\}_{j=1}^{M},\{G_{i}\}_{i=1}^{N})$triangleright Lift 2D features to 3D
21:$\{F_{i}\}_{i=1}^{N}\leftarrow\text{normalize}(\{F_{i}\}_{i=1}^{N})$triangleright Normalization
22:Return: 3D Gaussian-feature pairs $\{(G_{i},F_{i})\}_{i=1}^{N}$
4.2 Vision-Language 3DGS Pretraining
We first adapt the transformer encoder-decoder backbone from [51] to efficiently predict high-dimensional per-primitive features corresponding to collected 3DGS language labels. More specifically, our model $g(\cdot)$, parametrized by $\theta$, maps input Gaussians to their language features:
$\hat{F}=g_{\theta}(\{G_{i}\}_{i=1}^{N})\enspace,$ (1)
where $\hat{F}\in\mathbb{R}^{N\times d}$ is the predicted per-gaussian feature.
We apply three training objectives for supervision. Cosine similarity loss minimizes the angular difference between the predicted and ground truth language labels:
$\mathcal{L}_{\text{cos}}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\left(1-\frac{\hat{F}_{i}\cdot F_{i}}{||\hat{F}_{i}||\cdot||F_{i}||}\right),$ (2)
where $\mathcal{V}$ is the set of Gaussians with language feature labels. To enforce feature similarity in Euclidean space, we use L2 loss:
$\mathcal{L}_{2}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}||\hat{F}_{i}-F_{i}||^{2}\enspace.$ (3)
Lastly, we use an aggregated contrastive loss to encourage separation of class-level features. Instead of contrasting every Gaussian feature individually (which would be computationally prohibitive in large scenes), we apply class-wise mean pooling. For each semantic class $c$ with sufficiently many Gaussians, we randomly split its Gaussians into two disjoint sets $\mathcal{G}_{c}^{A}$ and $\mathcal{G}_{c}^{B}$, and compute the pooled features:
$\bar{F}_{c}^{A}=\text{Pool}(\hat{F},\mathcal{G}_{c}^{A})\,,\quad\bar{F}_{c}^{B}=\text{Pool}(\hat{F},\mathcal{G}_{c}^{B})\enspace.$ (4)
We then apply a bidirectional contrastive loss. First, we normalize $\bar{F}_{c}^{A}$ and $\bar{F}_{c}^{B}$ to unit length. Let $Z^{A}=\bar{F}^{A}(\bar{F}^{B})^{\top}/\tau$ and $Z^{B}=\bar{F}^{B}(\bar{F}^{A})^{\top}/\tau$, where each row in $\bar{F}^{A}$ or $\bar{F}^{B}$ corresponds to a different semantic class. The diagonal elements of $Z^{A}$ and $Z^{B}$ are the positive matches. Hence we compute a cross-entropy loss in both directions:
$\mathcal{L}_{\mathrm{contrast}}=\frac{1}{2|C|}\sum_{X\in\{A,B\}}\sum_{i\in C}-\log\frac{\exp(Z^{X}_{i,i})}{\sum_{j\in C}\exp(Z^{X}_{i,j})}\enspace,$ (5)
where $C$ is the set of semantic classes with sufficient Gaussians, $\tau$ is a learnable temperature controlling the softness of the distribution.
The total loss is the weighted sum $\mathcal{L}_{\text{total}}=\lambda_{\text{cos}}\mathcal{L}_{\text{cos}}+\lambda_{\text{L2}}\mathcal{L}_{\text{2}}+\lambda_{\text{con}}\mathcal{L}_{\text{contrast}}$, where $\lambda_{(\cdot)}$ denotes each weight. In practice, we found that applying the contrastive loss later during the training (i.e., “warm starting” with $\mathcal{L}_{\text{cos}}$ and $\mathcal{L}_{\text{2}}$) helps promote early feature learning while effectively refining class distinctions in later stages.
Through this training, our model learns to predict semantically rich language features for each Gaussian primitive, enabling downstream open-vocabulary scene understanding task without requiring additional finetuning or 2D input.
4.3 Self Supervised Pretraining
The proposed GaussianSSL method is presented in Fig. 2 (right). It incorporates multiple losses with different objective into SceneSplat’s large-scale pretraining.
Masked Gaussian Modeling. This part supervises the model to predict masked Gaussian primitives. Given a 3D Gaussian splatting scene represented by $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, where $G_{i}\in\mathbb{R}^{59}$, the training proceeds as follows. (1) A subset $\{G_{j}\}_{j=1}^{N^{\prime}}$ is sampled from $\mathcal{G}$ using a dense grid sampling of size $S$. (2) Samples are projected into a latent space with the embedding function $P$, in order to obtain tokens $E=P\left(\{G_{j}\}_{j=1}^{N^{\prime}}\right)\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$, where $d_{\text{e}}$ is the embedding dimension. (3) The tokens $E$ are masked with ratio $r\in[0,1]$, by replacing $N^{\prime}\cdot r$ randomly chosen tokens with a learnable mask token $t\in\mathbb{R}^{d_{\text{e}}}$, to obtain $T_{\text{m}}\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$. (4) The masked tokens $T_{\text{m}}$ are processed with the 3D backbone $g_{\theta}(\cdot)$ to obtain $\hat{T_{\text{m}}}=h_{\phi}(f_{\varphi}(T_{m}))$, where $f_{\varphi}(\cdot)$ is the encoder and $h_{\varphi}(\cdot)$ is the decoder. (5) The output tokens $\hat{T_{\text{m}}}$ are mapped to the input Gaussian space, with the reconstruction projector $\hat{G}_{m}={\Phi}(\hat{T}_{m})\in\mathbb{R}^{N^{\prime}\times F}$. (6) Finally, the $\mathcal{L}_{2}$ reconstruction loss is used between the predicted masked Gaussians and the original ones: $\mathcal{L}_{\text{MGM}}=\mathbb{E}_{G_{j}\sim\mathcal{G}}\left[\|G_{m}-\hat{G}_{m}\|_{2}^{2}\right]$.
Self-Distillation Representation Learning. Self-Distillation [31] learns augmentation-invariant representations by aligning predictions of a student network $\theta_{s}$ with an EMA-updated teacher $\theta_{t}$ [3]. For a batch of Gaussian scenes $\{\mathcal{G}_{n}\}_{n=1}^{B}$ (global/local views $G_{g}^{b}$, $G_{l}^{b}$), we extract tokenized bottleneck features $z\in\mathbb{R}^{M\times d_{e}}$, compute global representations $\bar{z}$ via mean pooling, and align student-teacher outputs via cosine similarity loss $\mathcal{L}_{\text{sim}}$ [52]. Feature diversity is regularized with a coding rate term $\mathcal{L}_{\text{cr}}$ [52, 23] resulting in $\mathcal{L}_{\text{DINO}}=\omega_{\text{sim}}\mathcal{L}_{\text{sim}}+\omega_{\text{cr}}\mathcal{L}_{\text{cr}}$ with corresponding weights. Inspired by [68, 52], the student network also predicts masked feature aligned with the teacher via an $\mathcal{L}_{\text{iBOT}}$, computed using cosine similarity. We propose to mitigate the decoder collapse issues by multi-task reconstruction $\mathcal{L}_{\text{MGM}}$, as coding rate regularization stabilizes only the hierarchical encoder. Following [11, 53], reconstruction is limited to weakly augmented views to avoid degradation.
Language-Gaussian Alignment. As shown in Sec. 4.2, the precomputed language feature enables effective knowledge distillation. For scenes with existing language labels, we seek to leverage them to further regulate self-supervised learning. However, the high dimensionality of these language features (dimension $N\times d_{L}$ ) can substantially increase the computational cost for supervision. To address this, we replace the original features with a compressed representation learned via an autoencoder[19], drastically reducing memory overhead while preserving semantic information. Similar to $\mathcal{L}_{\text{MGM}}$, we use $\mathcal{L}_{\mathrm{LA}}$ follow Eqs. 2 and 3 to train the network to predict low-dimensional language feature from unmasked neighbors.
5 Experiments
In this section, we evaluate the performance of vision-language pretraining on open-vocabulary task and examine the impact of large-scale Gaussian self-supervised pretraining on downstream indoor semantic segmentation. We further justify our design choices through ablation studies. Implementation details are provided in the supplement.
Method Training Source #Training Scenes ScanNet200 (200) Matterport3D (160) ScanNet++ (100)
f-mIoU f-mAcc f-mIoU f-mAcc f-mIoU f-mAcc
OpenScene [33] SN $\times$1 6.4 12.2 5.7 10.7 8.8 14.7
PLA [8] SN - 1.8 3.1 - - - -
RegionPLC [54] SN - 9.2 16.4 6.2 13.3 11.3 20.1
OV3D [15] SN - 8.7 - - - - -
Mosaic3D [21] SN - 13.0 24.5 8.6 17.8 16.2 27.1
SceneSplat (Ours) SN - 18.9 31.7 10.8 18.7 14.7 24.7
Mosaic3D [21] SN, SN++, ARKitS, MP3D, S3D $\times$24.3 15.7 28.3 13.1 27.7 18.0 29.0
SceneSplat (Ours) SN++ $\times$0.75 11.8 19.2 10.6 18.6 26.8 45.3
SceneSplat (Ours) SN, SN++ $\times$1.75 19.3 41.5 12.6 23.9 29.1 52.4
Table 2: Zero-Shot 3D Semantic Segmentation on the Fine-Grained ScanNet++ (100 classes) [57], Matterport3D (160 classes) [2] and ScanNet200 (200 classes) [5] Benchmarks. We report the foreground mean IoU (f-mIoU) and foreground mean accuracy (f-mAcc) excluding background classes (wall, floor, ceiling), following [54, 33, 8, 14]. denotes the official checkpoint and the results of the baselines are taken from [21]. Dataset abbreviations SN, SN++, ARKitS, MP3D, and S3D respectively denote ScanNet [5], ScanNet++ [57], ARKitScenes [1], Matterport3D [2] and Structured3D [66]. SceneSplat achieves noticeably better segmentation performance, i.e., a $5.9\%$ f-mIoU increase on the ScanNet200 benchmark when trained on a single source, and an $11.1\%$ f-mIoU increase on ScanNet++ when trained on two sources, while using significantly less training data compared to the concurrent work [21].
5.1 Vision-Language Pretraining
Tab. 2 reports the zero-shot 3D semantic segmentation results on the fine-grained ScanNet++ (100 classes) [57], Matterport3D (160 classes) [2] and ScanNet200 (200 classes) [5] benchmarks, where methods are trained on specified data sources. When trained on ScanNet, SceneSplat achieves the state-of-the-art results, leading to a $5.9\%$ and $2.2\%$ f-mIoU increase on the ScanNet200 and Matterport3D benchmarks respectively. By extending the training sources, we obtain a $3.6\%$ and $11.1\%$ f-mIoU increase on the ScanNet200 and ScanNet++ benchmark, falling behind only by $0.5\%$ f-mIoU compared to concurrent work on Matterport3D. Notably, [21] uses 14 times more training scenes to achieve its best results.
Fig. 3 shows the zero-shot segmentation results on evaluation scenes. SceneSplat not only achieves competitive segmentation performance but also correctly annotates the missing semantic labels (e.g., desks shown above). We demonstrate text-based queries on the inference results of predicted language features in Fig. 4. Our vision-language pretraining enables effective localization of complex objects within the scene.
Zero-Shot Prediction Ground Truth
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Qualitative Results of Zero-Shot 3D Semantic Segmentation on ScanNet++. SceneSplat demonstrates competitive zero-shot performance, note how our model correctly annotate the regions lacking ground truth labels, e.g., desks on the top row. Best viewed zoomed in and in color.
Figure 4: Text-Based 3DGS Scene Query. Given text queries and SceneSplat inference results for a 3DGS scene, we can effectively localize the corresponding splats (highlighted in red for queries ”Robot Arm”, ”Box”, and ”Keyboard”).
Method ScanNet20 (20) ScanNet200 (200) ScanNet++ (100)
mIoU mAcc mIoU mAcc mIoU mAcc
No-Pre 77.1 84.1 35.4 44.0 42.4 53.3
MGM 76.7 83.5 35.5 44.5 41.7 51.9
+DINO 77.0 84.6 35.9 46.1 42.0 52.4
+iBOT 77.2 84.2 35.2 44.3 41.1 52.0
+LA 77.2 84.2 34.7 44.4 41.4 52.5
Table 3: GaussianSSL Ablation Experiments. We adopt the pre-training on the Sceenssplat-7K dataset and report fine-tuning mIOU and mAcc on indoor semantic segmentation tasks. For details of specific losses please refer to Sec. 4.3 and Fig. 2.
Method ScanNet20 ScanNet200 ScanNet++
mIoU mAcc mIoU mAcc mIoU mAcc
PTv1 [65] 70.6 27.8
PTv2 [50] 75.4 30.2
PTv3 [51] 76.4 83.5 35.0 44.2 42.6 53.0
SceneSplat (Ours) 77.2 84.6 35.9 46.1 42.4 53.5
Table 4: Supervised Semantic Segmentation Experiments. We report our best results from Tab. 7 comparing against the state-of-the-art Point Transformer method.
5.2 Label-free 3DGS Pretraining
We conduct extensive pretraining experiments with the proposed GaussianSSL method. To assess the efficacy of the pretrained model, we report the segmentation results in Tab. 3. Our method achieves a +0.1% improvement over supervised-only baselines on ScanNet20 and +0.5% on ScanNet200, while observing a performance drop on ScanNet++ primarily due to pretraining dataset quality variations (Tab. 1). Furthermore, compared to our reproduced implementation of PTv3  [51], we outperform by +0.8% on ScanNet20 and +0.9% on ScanNet200 (Tab. 4). More qualitative results are in the supplement.
5.3 Further Statistical Evaluation
Method ScanNet200 (200) ScanNet++ (100)
f-mIoU f-mAcc f-mIoU f-mAcc
Language Labels 22.8 35.9 22.6 46.5
SceneSplat 18.9 31.7 26.8 45.3
Table 5: SceneSplat Inference Results vs. Collected Language Labels on Zero-Shot 3D Semantic Segmentation. The result on ScanNet++ shows the inference performance can be even better than using the collected labels. SceneSplat here is trained on the single dataset respectively.
SceneSplat Inference Results vs. Collected Language Labels. One may assume the collected language labels cap the upper bound zero-shot performance since they provide the supervision signal during vision-language pretraining. Interestingly, this is not always the case. Tab. 5 compares the performance of SceneSplat inference features with the collected language labels. On ScanNet++, ours outperforms the labels with a $4.2\%$ increase in f-mIoU. Although the collected labels are not perfect, large-scale pretraining is able to filter noise and learn meaningful patterns.
Figure 5: Distribution of SceneSplat Zero-Shot Semantic Segmentation mIoU w.r.t. Input 3DGS Scene PSNR. Reported on the Matterport3D test split labeled in 21 semantic classes, the box plot shows a clear positive trend between the input 3DGS scene training PSNR and the resulted mIoU once applied SceneSplat language pretraining for zero-shot semantic segmentation. This encourages the careful curation of the collected 3DGS scene dataset.
Impact of the Input 3DGS Scene PSNR on Open-Vocabulary Performance. Reported on the Matterport3D test split with 370 scenes, Fig. 5 indicates a clear positive trend of the input 3DGS scene PSNR of the training views and the resulted zero-shot mIoU performance with SceneSplat model. Low PSNRs usually come out of blurry input images, poor Gaussian centers optimization and insufficient scene coverage, where the 3DGS parameters cannot well resolve the scene. The trend here highlights the importance of data curation on the collected 3DGS scenes.
Figure 6: Overall and Class-Wise IoU Changes w.r.t. to the Nearest Neighbor Number During Majority Voting. We evaluate SceneSplat using different nearest 3DGS neighbors for zero-shot task at the point locations on Scannet++ validation split. Overall mIoU increases with different class-wise relative IoU changes.
SceneSplat Training Input ScanNet200 (200) ScanNet++ (100)
mIoU mAcc mIoU mAcc
Point Parameters 17.1 27.9 23.8 40.2
3DGS Parameters 18.9 31.7 26.8 45.3
Table 6: Zero-Shot Performance of Using Point Clouds vs. 3DGS for SceneSplat Vision-Language Pretraining. SceneSplat trained on 3DGS parameters consistently outperforms the variant trained on point cloud properties. The models here are trained on the single dataset respectively.
Contrastive Loss ScanNet200 (200) ScanNet++ (100)
f-mIoU f-mAcc f-mIoU f-mAcc
w/o 13.7 22.5 19.6 34.4
always apply 13.2 23.4 23.2 39.3
last 75% epochs 15.7 24.0 23.8 40.2
Table 7: Ablation on Contrastive Loss During Vision-Language Pretraining Using Subsets. Having a warm-up period and applying the contrastive loss later leads to better performance.
Method Steps Required Runtime / Scene
Occam’s LGS 2D fusion + lifting 107 min
SceneSplat single inference 0.24 min
Table 8: Runtime. SceneSplat is significantly faster compared to the current fastest language-embedded 3DGS method.
Nearest Neighbors Voting During Zero-Shot Experiments. The centers of input Gaussians differ from the point locations where the semantic predictions are evaluated on, thus we have to aggregate predictions from neighboring Gaussians. We carry out majority voting using the nearest neighboring Gaussians for each evaluating location. Fig. 6 ablates the number of nearest neighbors on the IoU results using the ScanNet++ validation split. We observe the overall trend of mIoU increase w.r.t. to the number of nearest neighbors and list the classes with top relative changes. To balance the performance and inference speed, 25 nearest neighbors are selected for voting.
Effectiveness of Using 3DGS in Vision-Language Pretraining Compared to Point Clouds. To further justify the effectiveness of using 3DGS parameters for scene understanding, we apply the same vision-language pretraining on point properties (color and normal). Tab. 6 indicates that the model takes point properties as input consistently get outperformed by SceneSplat using 3DGS parameters.
Ablation on Contrastive Loss in the Lang-Aware Pretraining. Tab. 7 ablates the contrastive loss applied during vision-language pretraining, where applying contrastive loss at the late stage of training outperforms other variants.
Runtime vs. Per-Scene Language Gaussian Splatting Method. Thanks to the feed-forward ability after vision-language pretraining, SceneSplat is shown in Tab. 8 to be 445.8 times faster than the state-of-the-art Language-embedded Gaussian Splatting method, as there is no need for 2D feature extraction and fusion.
6 Conclusion
In this work, we introduce SceneSplat, the first large-scale 3D scene understanding model for indoor environments operating directly on 3D Gaussian splats. Powered by SceneSplat-7K, a dataset comprising 6,868 scenes, we propose a novel 3D Gaussian splat encoder that generates semantic features in a single pass, enabling open-vocabulary scene recognition without relying on 2D fusion. Through self-supervised techniques, we unlock label-free 3DGS pretraining at the scene level. Our approach achieves state-of-the-art performance in zero-shot semantic segmentation, establishing new benchmarks and laying the foundation for future advancements in open-vocabulary 3D understanding.
[…]
Supplementary Material
A Implementation Details
A.1 Language Label Collection
Many existing methods align 3D primitives with text embeddings from vision-language models (VLM) [8, 21]. While effective for basic understanding, this approach inherently limits the information captured, as textual descriptions typically only convey categorical and spatial properties, lacking fine-grained visual details and relationships. Methods employing visual captioning models [26, 59] face similar challenges, as they struggle to describe all aspects of the target scene content. Even detailed captions inevitably result in information loss during the text embedding alignment process. Differently, we directly align Gaussians with the image embedding space of VLM, thereby preserving richer latent semantic information.
Dynamic Weighting Mechanism. Unlike previous approaches that use a single tight crop around each segment, we adapt the three-crop strategy [49] with dynamic weighting to capture the context. For each segment identified by SAMv2, we extract three distinct features: (1) Global feature $f_{g}$ from the entire RGB frame, capturing the full scene context; (2) Local feature $f_{l}$ from the image crop with background; (3) Masked feature $f_{m}$ from the image crop without background, focusing solely on the object. We use SAMv2/sam2-hiera-large and SigLIP2/siglip2-base-patch16-512 models.
Our dynamic weighting mechanism combines the three extracted features through the following equations. First, we calculate the cosine similarity between features with and without background and use it to create a fused local feature:
$r_{lm}$ $=\text{sim}(f_{l},f_{m})$ (6)
$F_{l}$ $=r_{lm}\cdot f_{m}+(1-r_{lm})\cdot f_{l},$ (7)
$F_{l}$ $=\text{normalize}(F_{l})$ (8)
We then compute the similarity between this fused local feature and the global feature to determine the final weights:
$\phi_{lG}$ $=\text{sim}(F_{l},f_{g}),\quad w_{i}=\text{softmax}(\phi_{lG})$ (9)
$w_{g}$ $=w_{i},\quad w_{m}=(1-w_{i})\cdot r_{lm},$ (10)
$w_{l}$ $=(1-w_{i})\cdot(1-r_{lm})$ (11)
The final fused feature is computed as a weighted combination and normalized:
$f_{s}$ $=w_{g}\cdot f_{g}+w_{l}\cdot f_{l}+w_{m}\cdot f_{m},\quad f_{s}=\text{normalize}(f_{s})$ (12)
This mechanism adaptively balances global context influence via $w_{g}$, local context with background via $w_{l}$, and object-specific features via $w_{m}$. These three features are dynamically combined to create a representation that adapts based on the segment’s relationship with its context. For objects highly integrated with their surroundings (e.g., a keyboard in front of a monitor), background-inclusive features receive a higher weight. For isolated objects (e.g., a coffee mug), background-excluded features dominate.
A.2 Vision-Language Pretraining
Our vision-language 3DGS pretraining model is built on a transformer encoder-decoder architecture adapted from [51]. As detailed in Tab. F, the encoder consists of 4 stages with depths [2,2,2,6], channels [32,64,128,256], and heads [2,4,8,16], while the decoder employs 3 stages with depths [2,2,2], channels [768,512,256], and 16 attention heads each. We process 3D Gaussian primitives with all parameters (center, color, opacity, quaternion, and scale). The model is trained with AdamW optimizer (initial LR = 0.006, weight decay = 0.05) using a OneCycle scheduler with cosine annealing. Our loss weights are set to $\lambda_{\text{cos}}=1.0$, $\lambda_{\text{L2}}=1.0$, and $\lambda_{\text{con}}=0.02$, with contrastive loss (temperature $\tau=0.2$) activated only in the later $75\%$ of training. During training, we employ extensive data augmentation including random rotations, scaling, flipping, jittering, and elastic distortions (see Tab. H). We train for 800 data epochs using 4 NVIDIA H100 GPUs.
A.3 Masked Gaussian Modeling
For all self-supervised pretraining experiments on the full dataset, we employ 4 NVIDIA H100 GPUs (80GB each) and the AdamW optimizer, with learning rates of $1\times 10^{-3}$ for the embedding layer and $1\times 10^{-4}$ for the attention blocks, coupled with a weight decay of $1\times 10^{-3}$. We use mixed-precision training (FP16) to accelerate convergence and reduce memory overhead. In total we train for 300k steps. We use an $\mathcal{L}_{2}$ loss to enforce consistency between predicted and ground truth Gaussian parameters, focusing only on masked regions. For parameter reconstruction, we design a three-layer MLP projector for each Gaussian attribute, incorporating output activations that adhere to physical constraints: Color use a $\tanh$ activation (bounded in $[-1,1]$), Opacity and Scale use a sigmoid activation(bounded in $[0,1]$) because in most indoor scenes, most Gaussians have small scales, Rotation, and Normals applies $\tanh$ followed by $\ell_{2}$-normalization to ensure valid quaternions and normals.
A.4 Self-Distillation Representation Learning
In self-distilled representation learning, we adopt the framework of [52], combining DINO loss and coding rate regularization on pooled encoder features. We configure a batch size of 24 and set the grid sampling resolution to 0.02 for partitioning Gaussian scenes. After grid sampling, scenes are randomly partitioned into base views $\mathcal{G}_{b}$, which are then used to generate global and local crops. For global crops, we randomly select a Gaussian splat from $\mathcal{G}_{b}$ and sample its $K$ neighbors, where $K\sim[0.4,1.0]\times 256,000$. For local crops, we sample $K\sim[0.1,0.4]\times 102,400$. Each batch includes $N_{g}=2$ global views and $N_{l}=3$ local views to balance scene coverage and granularity. For the DINO loss $\mathcal{L}_{\text{DINO}}$, we extract tokenized features from the student encoder $f_{\theta}(\cdot)$ and teacher encoder $f_{\phi}(\cdot)$, where $\phi$ is updated via exponential moving average (EMA). A global representation $\bar{z}$ is obtained by mean pooling over spatial tokens. This representation is projected into a latent space using a 3-layer MLP $P_{\text{DINO}}$ (dimensions: 2048 → 2048 → 256) followed by $\ell_{2}$-normalization. Then we calculate similarity loss:
$\mathcal{L}_{\text{sim}}=\frac{1}{N_{g}\times N_{l}}\sum_{i=1}^{N_{l}}\sum_{j=1}^{N_{g}}\frac{P_{\text{DINO}}(\bar{z_{l}}^{(i)})\cdot P_{\text{DINO}}(\bar{z_{g}}^{(j)})}{\|P_{\text{DINO}}(\bar{z_{l}}^{(i)})\|\|P_{\text{DINO}}(\bar{z_{g}}^{(j)})\|}$ (13)
For regularization, we use the negative of coding rate:
$R_{\epsilon}(\Gamma):=\frac{1}{2}\text{logdet}\left(\mathbf{I}+\frac{d}{\epsilon^{2}}\Gamma\right),$ (14)
$R_{\epsilon}$ approximates the rate-distortion function of a Gaussian random variable with covariance $\Gamma$, becoming exact as $\epsilon\to 0$. More specifically, it quantifies the spread of the covariance, even when the underlying variables are not strictly Gaussian.
For the iBOT loss, we mask 50% of global views with masking ratios sampled uniformly from $r\sim[0.2,0.7]$. Masked tokens are replaced by a learnable token ($T_{\text{mask}}$). Note here the masked global view embedding feature goes through the student network and outputs $\hat{T^{\prime}}(\theta_{s})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$ while $d_{\text{r}}$ note the representation dimension. and the unmasked global view embedding forward to teacher network and output $\hat{T^{\prime}}(\theta_{t})\in\mathbb{R}^{N^{\prime}\times d_{\text{r}}}$. We detach the teacher’s output and calculate the simple iBOT loss using 3-layer MLP $P_{\text{iBOT}}(\cdot)$ (dimensions: 256 → 256 → 32) as projector:
$\mathcal{L}_{\text{iBOT}}=\frac{1}{|M|}\sum_{i\in M}\underbrace{\frac{P_{\text{iBOT}}(\hat{T^{\prime}}(\theta_{s})^{(i)})\cdot P_{\text{iBOT}}(\hat{T^{\prime}}(\theta_{t})^{(i)})}{\left\|P_{\text{iBOT}}(\hat{T^{\prime}}(\theta_{s})^{(i)})\right\|\left\|P_{\text{iBOT}}(\hat{T^{\prime}}(\theta_{t})^{(i)})\right\|}}_{\text{Pairwise Cosine Similarity}}$ (15)
Only the masked regions $M$ contribute to the iBOT loss calculation, Following [52], we simplify the original iBOT implementation by removing the centering operation and online tokenizer, replacing them with a pairwise cosine-similarity loss for feature alignment. When MGM loss is enabled then for one global view we will employ less aggressive augmentation and this view will contribute to both iBOT loss and MGM loss.
A.5 Autoencoder for feature compression
Building on the idea from [36], we train a scene-specific language autoencoder on precomputed vision-language embeddings. This model compresses high-dimensional SigLIP features into a low-dimensional latent space, enabling efficient storage of language features within Gaussian representations. The encoder (by default 5-layer architecture: [384, 192, 96, 48, 16]) generates compact latent codes, while the symmetric decoder (by default 5-layer architecture: [48, 96, 192, 384, 768]) reconstructs the original high-dimensional embeddings. This design reduces memory overhead while preserving semantic fidelity. Unlike [36], our model is trained on 3D Gaussian features rather than 2D image-level data. For detailed ablations on the autoencoder architecture and training paradigm, see Tab. B.
A.6 Gaussian Language Alignment.
We follow the same loss functions in Eqs. 2 and 3 for feature alignment. When combining MGM with language alignment, we adhere to the MGM process by masking input tokens and applying the compressed language alignment loss only to the masked regions. Given the comparable magnitudes of all loss terms, we assign equal weights ($\omega_{\text{MGM}}=\omega_{\text{DINO}}=\omega_{\text{iBOT}}=\omega_{\text{LA}}=1.0$) to maintain the balance of different objective during training.
B Further Experimental Results
B.1 SceneSplat Zero-shot Segmentation
Figure A: Comparison of SceneSplat Zero-Shot Versus Fully Supervised Segmentation Across Object Classes. Notably, our zero-shot segmentation after vision language pretraining demonstrates better performance for 18 object classes, predominantly small objects such as headphones, printer, pot, and clothes hanger.
Zero-shot vs. Fully Supervised. Fig. A compares the class-wise zero-shot 3D segmentation results with those of the current state-of-the-art supervised method [51]. Points above the diagonal line (colored red) represent classes where zero-shot segmentation outperforms fully supervised approaches, while points below (colored blue) indicate the opposite. Our zero-shot segmentation achieves superior performance for 18 object classes in the ScanNet++ benchmark, predominantly small objects such as headphones, printers, pots, and clothes hangers. These results demonstrate the robust prior knowledge acquired by our model through vision-language pretraining.
More Qualitative Zero-shot Segmentation Results. Fig. C presents more qualitative zero-shot semantic segmentation results on ScanNet++ validation scenes. SceneSplat effectively segments the scenes and helps annotate regions with missing labels in the ground truth.
Consistency Issue in Label Collection. Tab. A presents our zero-shot segmentation results on the ScanNet20 benchmark. We observe that performance on this particular benchmark is significantly lower compared to the state-of-the-art results we achieve on the other three benchmarks. Through detailed analysis, we identify that the issue stems from inconsistencies in the 3DGS language label collection process. Specifically, the SAMv2+SigLIP2 pipeline we employ does not guarantee temporal consistency, especially for large background objects such as walls and floors, as illustrated in Fig. B. This inconsistency results in corrupted feature fields for Gaussians associated with these regions, subsequently leading to reduced zero-shot segmentation performance. Addressing this temporal consistency issue remains an important direction for our future work.
Method Training Source ScanNet20 (20)
f-mIoU f-mAcc
OpenScene ScanNet 57.5 72.4
Mosaic3D ScanNet 65.0 82.5
PLA ScanNet 19.1 41.5
RegionPLC ScanNet 55.6 76.3
OV3D ScanNet 64.0 76.3
SceneSplat ScanNet 35.4 57.9
Table A: Zero-Shot 3D Semantic Segmentation on ScanNet20 Benchmark. The results for the baselines are taken from [21]. Ours observes many faulty predictions and have poor performance, we identify the issue in the inconsistency during 2D feature map collection when labeling Gaussians, which leads to corrupted 3DGS-feature pairs.
Refer to caption Refer to caption
Figure B: Consistency Issue During 2D Vision-Language Feature Map Collection on ScanNet. Erroneous 2D feature maps are collected shown by the mislabeled regions in the neighboring figures. The root cause is that the SAMv2+SigLip2 process we use does not guarantee temporal consistency.
Figure C: Qualitative Zero-shot Semantic Segmentation Results on ScanNet++ Validation Scenes. SceneSplat effectively segments the scenes and helps annotate regions with missing labels in the ground truth.
B.2 Gaussian Self-Supervised Pretraining
Method Scannet++ 2D Scannet++ 3D
L2 cosine L2 cosine
VAE16 layer5 0.008 0.182 0.005 0.028
VAE16 layer6 0.008 0.172 0.005 0.023
VAE64 layer5 0.007 0.081 0.004 0.014
VAE64 layer6 0.007 0.079 0.004 0.013
Table B: AutoEncoder Ablation Experiments. We report the feature compression performance with $\mathcal{L}_{2}$ loss and cosine similarity on unseen 3D language label. We ablate on different autoencoder architectures and training sources.
In Tab. B, we present an ablation study comparing various autoencoder architectures trained on the ScanNet++ training set and evaluated on the validation set. In the “Scannet++ 2D” columns, the autoencoder is trained using SigLIP2 features preprocessed from images, while in the “Scannet++ 3D” columns, it is trained directly on Gaussian SigLIP2 features. At inference, we measure both the L2 distance and cosine similarity on 3D Gaussian SigLIP2 features. The results indicate that training the autoencoder directly on 3D data yields notably better performance, whereas increasing the network depth from layer5 to layer6 leads to only marginal improvements. Moreover, a 64-dimensional latent space consistently outperforms the 16-dimensional counterpart. Consequently, and as detailed in Sec. A.5, we adopt the 3D-trained autoencoder with layer5 and a 16-dimensional latent space as our default configuration, striking a favorable balance between efficiency and accuracy.
Method ScanNet20 (20)
mIoU mAcc
LA16 76.3 83.9
LA64 75.8 82.2
LA16 (MGM) 76.2 83.7
Table C: Language Alignment Loss Ablation. We conduct an ablation study on low-dimensional language feature compression and Masked Gaussian Modeling (MGM)-based pretraining to evaluate their impact on semantic segmentation performance.
We evaluate the impact of language feature alignment loss dimensionality (64 vs.16) when pretraining exclusively on ScanNet++ and testing on the ScanNet20 benchmark. As shown in Tab. C, We implement two 3-layer MLP architectures: one with uniform dimensions (64 → 64 → 64) for 64 dimension language feature and another with progressively reduced dimensions (64 → 32 → 16). We observe that reducing the latent dimension from 64D to 16D improves mIoU by +$0.5\%$ in mIoU and +0.7% in mAcc. This suggests that lower-dimensional language features with hierarchical dimensionality reduction preserve critical indoor scene semantics information.
We visualize pretraining results using Masked Gaussian Modeling (MGM) and language feature alignment loss in
Figure D: Self-supervised Reconstructions across Multiple Scenes. Each row shows (left to right) the unmasked input, masked scene, reconstruction, and a PCA projection of features
. The model effectively reconstructs masked Gaussian parameters (e.g., position, scale, rotation) and predicts semantically meaningful language-aligned features for indoor scene understanding.
Tab. D evaluates how masking ratio and grid size in Masked Gaussian Modeling (MGM) affect downstream semantic segmentation performance. Key observations include: (1) Small Masking ratios (0.4) degrade mIoU by $0.9\%$ due to insufficient context learning compared with the 0.6. Using a larger grid size (0.15m) for masking reduces fine-grained detail recovery, leading to a 0.5% drop in mIoU compared to a 0.05m grid size.
Mask Ratio ScanNet20 Mask ScanNet20
mIoU mAcc Size mIoU mAcc
0.4 76.1 83.6 0.02 76.9 84.5
0.5 76.4 84.1 0.05 77.0 84.5
0.6 77.0 84.4 0.10 76.6 83.8
0.7 76.7 83.8 0.15 76.5 84.1
Table D: Masked Gaussian Modeling Ablation Experiment. We analyze the impact of masking ratios and grid sizes in Masked Gaussian Modeling (MGM) on semantic segmentation performance using the ScanNet20 benchmark.
Method ModelNet10 (10) Omniobject3D (83)
Linear MLP-3 Linear MLP-3
MGM 77.3 83.1 50.3 57.5
+DINO 74.5 80.2 40.5 42.4
+iBOT 65.2 73.8 38.4 40.8
+LA 68.1 74.8 26.5 31.2
MGM+LA 83.1 84.6 47.3 53.2
Table E: Cross Domain Linear Probe Experiments. We report the mAcc results in the ablation of GaussianSSL on object-level classification tasks using the linear probe.
To evaluate global scene understanding, we design a cross-domain classification task (Tab. E) using object-level Gaussian splats from [28, 29]. We freeze the encoder and embedding layers, training only (1) A linear probe (final layer) to map features to logits. (2) A 3-layer MLP (512 → 256 → classes) for nonlinear evaluation. We find that adding progressively the DINO, iBOT, and language alignment (LA) losses degrade classification accuracy, By using Masked pretraining and language alignment loss achieves the best performance on ModelNet10 (furniture), while MGM alone excels on OmniObject3D (common objects). In Furniture datasets (e.g., ModelNet10) achieve the best 80% accuracy, while challenging object datasets (e.g., OmniObject3D) peak at  50%. The misalignment with scene-level benchmarks stems from domain gaps: the model maps architecturally distinct objects (e.g., fridge, oven) to similar scene-level semantics (e.g., ”kitchen”).
Figure E: Comparison of Scene Query Results Using Our Predictions and GT Language Labels on ScanNet++.
C Datasets Curation and Statistics
C.1 Analysis
Gaussian Splatting (GS) has demonstrated strong reconstruction capabilities, but its performance is heavily dependent on raw dataset quality and input conditions. We analyzed both quantitative metrics and visualization results of reconstructed scenes and identified several common failure cases. The primary issues include holes, floating artifacts, and non-smooth surfaces. Through our analysis, we categorize the root causes into four main factors.
Lack of Frames and Filming Angles. Scenes with fewer than 400 RGB frames often fail to capture a complete representation of the environment. Limited filming angles result in missing perspectives, leading to incomplete indoor structures. Fig. F from the ARKitScenes dataset show partial room reconstructions where corners or ceiling details are lost.
Figure F: Example Incomplete Scenes in ARKitScenes.
Motion Blur and Camera Instability. As seen in some blurry input examples Fig. G in the 3RScan dataset, the rapid, unsteady filming leads to edge deformation, ghosting, and smeared textures, and in unrealistic reconstructions, the sharp edges are lost, significantly reducing the fine details in the rendered GS.
Figure G: Blurry Scenes from 3RScan.
Indoor-outdoor Lighting Changes. Scenes with dynamic indoor-outdoor lighting changes, such as the case from the Hypersim dataset, exhibit severe blurring and loss of surface textures. Large glass ceilings, skylights, and reflective floors further disrupt GS training, as the algorithm struggles with light inconsistency and high contrast between illuminated and shadowed areas.
Challenges with Transparent and Glass Objects. Transparent and glass objects pose a unique challenge, as GS often fails to accurately capture windows or furniture glass, resulting in missing elements or floating artifacts. This issue arises from the inherent difficulty in rendering transparency, where reflection and refraction introduce complexity beyond the current GS capabilities.
C.2 Visualization
The SceneSplat-7K dataset achieves an impressive average PSNR of 28.17 dB through all the different sources. We provide visualizations that showcase the photorealistic appearance rendering. See Sec. C.2.
ScanNet++ GS. Ground truth (left) and 3DGS rendering results (right).
Hypersim GS. Ground truth (left) and 3DGS rendering results (right).
Replica Office0 GS. Ground truth (left) and 3DGS rendering results (right). PSNR: 45.55 dB.
Replica Room2 GS. Ground truth (left) and 3DGS rendering results (right). PSNR: 41.57 dB.
Figure L: 3RScan GS. Ground truth (left) and 3DGS rendering results (right). PSNR: 34.43 dB.
D Architecture Designs
Config Value
embedding depth 2
embedding channels 32
encoder depth [2, 2, 2, 6]
encoder channels [32, 64, 128, 256]
encoder num heads [2, 4, 8, 16]
encoder patch size [1024, 1024, 1024, 1024]
decoder depth [2, 2, 2]
decoder channels [768, 512, 256]
decoder num heads [16, 16, 16]
decoder patch size [1024, 1024, 1024]
down stride [$\times$2, $\times$2, $\times$2, $\times$2]
mlp ratio 4
qkv bias True
drop path 0.3
Table F: Model Configs for Vision-Language Pretraining.
Config Value
embedding depth 2
embedding channels 32
encoder depth [2, 2, 2, 6, 2]
encoder channels [32, 64, 128, 256, 512]
encoder num heads [2, 4, 8, 16, 32]
encoder patch size [1024, 1024, 1024, 1024]
decoder depth [2, 2, 2, 2]
decoder channels [64, 64, 128, 256]
decoder num heads [4 4, 8, 16]
decoder patch size [1024, 1024, 1024]
down stride [$\times$2, $\times$2, $\times$2, $\times$2]
mlp ratio 4
qkv bias True
drop path 0.3
Table G: Model Configs for GaussianSSL Pretraining and Downstream Semantic Segmentation.
Augmentations Parameters Global View Local View
Base Transform
random rotate axis: z, angle: [-1, 1], p: 0.5
axis: x, angle: [-1 / 64, 1 / 64], p: 0.5
axis: y, angle: [-1 / 64, 1 / 64], p: 0.5
random scale scale: [0.9, 1.1]
random flip p: 0.5
random jitter sigma: 0.005, clip: 0.02
elastic distort params: [[0.9, 0.1]]
grid sampling grid size 0.02
Global Base Transform -
random flip p: 0.5
random crop ratio: (0.4, 1.0) max: 256000
Global Transform 0 -
random color jitter b:0,4, c:0.4, s:0.2 hue:0.1 p:0.8
random grayscale p: 0.2
random Gaussian blur p: 1.0
Global Transform 1 -
random dropout dropout ratio: 0.2, p: 0.2
random color jitter b:0,4, c:0.4, s:0.2 hue:0.1 p:0.8
random grayscale p: 0.2
random Gaussian blur p: 0.2
Local Base Transform -
elastic distort params: [[0.2, 0.4], [0.8, 1.6]]
random flip p: 0.5
random crop ratio: (0.1, 0.4) max: 256000
Local Transform -
random dropout dropout ratio: 0.2, p: 0.2
random color jitter b:0,4, c:0.4, s:0.2 hue:0.1 p:0.8
random grayscale p: 0.2
random Gaussian blur p: 0.5
Table H: Data Augmentations. Following [31], we design the data augmentation for 3D scenes for global and local views.
We detail our network design choices in Tabs. F and G. For the 3D backbone, we adopt the state-of-the-art Point Transformer [51], enhanced with Flash Attention to substantially improve computational efficiency. To optimize feature embedding for large scale Gaussians, we use sparse convolutions, which preserve geometric details while minimizing memory overhead.
E Dataset License
SceneSplat-7K is constructed from several established indoor datasets, each attached with a specific license: ARKitScenes [1] (Apple Open Source License), Replica [45] (Replica Research License), ScanNet [5] (ScanNet Terms of Use), ScanNet++ [57] (ScanNet++ Terms of Use), Hypersim [42] (CC BY-SA 3.0), 3RScan [48] (CC BY 4.0), and Matterport3D [2] (non-commercial license). For sources that permit redistribution, we plan to release our data on Hugging Face under their respective terms of use. For datasets that require special permission, we will co-host the data only after receiving approval from the original teams.
F Limitations
Our dataset quality is largely influenced by the source indoor datasets. Low resolution and blurry images can cause floating artifacts. Downstream tasks are also limited by the original dataset’s annotations. Temporal inconsistency in the current language label obtaining process needs to be addressed as it can pollute the vision-language pretraining. Also, we plan to add the bounding boxes [13] and language descriptions [27] as next step.
G Impact Statement
This work introduces SceneSplat-7K, the first large-scale dataset of 3D Gaussian Splatting (3DGS) specifically designed for indoor scene understanding. By providing over 7,000 annotated scenes, it enables standardized benchmarking for 3DGS-based reasoning in vision tasks, paving the way for more robust 3D scene comprehension. The proposed framework for vision-language pretraining tailored to 3DGS enhances semantic alignment by incorporating language-annotated 3D Gaussian splats, which establish clear connections between latent representations and 3D scenes. Leveraging knowledge distillation from 2D foundation models, our approach pushes the boundaries of 3D vision-language understanding, achieving state-of-the-art performance in open-vocabulary tasks. Furthermore, our model, combining contrastive learning and masked point modeling (MPM), significantly outperforms existing methods in 3D semantic segmentation. This work lays a critical foundation for advancing scalable 3D scene understanding, with broad implications for autonomous systems, robotics, and augmented reality applications. By releasing the dataset, model, and code publicly, we foster further innovation and facilitate the development of generalizable solutions for 3D vision tasks.